# Expanded Generalized Needlet Internal Linear Combination (eGNILC) Framework for the 21-cm Foreground Removal

### v1

@@ Abstract
Intensity mapping (IM) of the neutral hydrogen 21-cm line has been studied as a promising technique to probe the large-scale structure (LSS) of the Universe (Madau et al., 1997; Battye et al., 2004; Masui et al., 2013; Santos et al., 2005; Bull et al., 2015; Santos et al., 2015; Bigot-Sazy et al., 2015, 2016; Li & Ma, 2017; Harper et al., 2018; Bacon et al., 2018; Xu et al., 2018; Anderson et al., 2018). Instead of measuring the emission from individual galaxies, IM measures integrated emission from a voxel of many galaxies in which the total signal can be detected by radio telescopes. Therefore, the integrated Hi  flux density in each pixel of sky maps contains multiple galaxies’ contributions to the signal.
Many experiments have been conducted or proposed to detect the Hi  power spectrum from auto- or cross-correlations. The techniques are mainly two kinds, the interferometer array and the single-dish (SD) measurement. The interferometer array measures the Fourier transform of the sky brightness in $uv$-plane (e.g. visibilities), which directly corresponds to the power spectra of the 21-cm signal. These include the Low-Frequency Array (LOFAR; Gehlot et al. 2019; Mertens et al. 2020), the Precision Array for Probing the Epoch of Reionization (PAPER; Parsons et al. 2010; Gogo et al. 2022), the Murchison Widefield Array (MWA; Trott et al. 2020), and the Hydrogen Epoch Reionization Array (HERA; DeBoer et al. 2017; Abdurashidova et al. 2022a, b; The HERA Collaboration et al. 2022). The single-dish experiment, on the other hand, measures the specific flux of the 21-cm signal by direct imaging, which has achieved several detections in recent years via cross-correlation, including the HIPASS cross-correlation with 6dFGS galaxies (Pen et al., 2009), Green Bank Telescope (GBT) cross-correlation with DEEP2 galaxy survey (Chang et al., 2010) and with WiggleZ survey (Masui et al., 2013), Parkes telescope cross-correlation with 2dFGS galaxies (Anderson et al., 2018), and stacking of Parkes 21-cm fields with galaxy halos (Tramonte & Ma, 2020) and filaments (Tramonte et al., 2019). More recently, Cunnington et al. (2022) utilized $10.5$-hour observations of MeerKAT ($64$-dishes) drifted scan over $\sim 200\deg^{2}$ and cross-correlated it with WiggleZ galaxies, and achieved $7.7\sigma$ detection of the cross-correlation signal. In the near future, Five-hundred-meter Aperture Spherical radio Telescope (FAST, Nan et al. 2011; Bigot-Sazy et al. 2016), BINGO (Battye et al., 2012, 2013; Dickinson, 2014; Battye et al., 2016; Wuensche, 2019; Yohana et al., 2019; Abdalla et al., 2022; Wuensche et al., 2022), and SKA-MID (Bacon et al., 2018) will also utilize the SD IM technique to measure the 21-cm brightness of the sky and probe the LSS.
@@
This paper is organized as follows. In Sec. 2, we describe the models for the Hi  signal, foreground components, and instrumental noise. In Sec. 3, we present the mathematical framework of the eGNILC and make an extensive discussion of its function on foreground cleaning. Then in Sec. 4, we demonstrate the performance of the foreground removal methods, and apply it to the simulated data of the BINGO and SKA-MID configurations in Sec. 5. The conclusion will be presented in the last section.
@@ 2.1 Hi  21-cm intensity map
where in the last line, we utilized the fact that the Hi  signal traces the underlying matter density field on large scales after reionization. $\delta^{(3)}_{\mathrm{D}}$ is the 3-D Dirac delta function, $P_{\textsc{Hi}}\,$ and $P_{\mathrm{m}}$ are the Hi  and matter power spectra, $D(z)$ is the growth factor, and $b$ is the bias between the spatial distribution of the Hi  and the matter field. Then by ensemble-averaging the quadratic $a_{\ell m}$ (Eq. (6)), we obtain the angular power spectrum of 21-cm fluctuations in redshift bins ($z,z^{\prime}$) (see also Battye et al. 2013)
@@ 2.2.2 Unresolved random sources
The spatial patterns are different in the DCT of power-law foregrounds because the synchrotron map is spatially smooth while temperature fluctuations are more dramatic in the free-free and point source maps. The unresolved maps of both synchrotron and point sources are stochastic “noises”. Because the angular power spectrum of the unresolved synchrotron decreases faster ($\alpha$ is larger) than that of the unresolved point-source map, the map of the unresolved synchrotron foreground looks smoother than the unresolved point-source map. The horizontal stripes of the DCT images also suggest such spatial smoothness, while descending mode coefficients reflect the correlation of the $\nu$ and $\nu^{\prime}$ angular power spectrum. An outstanding feature of the Hi  signal is that the mode coefficients do not decay visibly.
@@ 2.2.3 Instrumental noise
The most common instrumental noise is the Gaussian distributed thermal noise. The rms of pixel noise is related to the system temperature, integration time and frequency resolution (Bigot-Sazy et al., 2015; Olivari et al., 2016)
@@
where $f_{\rm k}$ is the knee frequency, $\alpha$ is the spectral index, $\omega$ is the inverse spectroscopic frequency wavenumber, $\Delta\nu$ is the total receiver bandwidth, $\beta$ is used to parametrize the spectral index of the PSD [^3] We remind the reader not to confuse the meaning of $\alpha$ and $\beta$ here with the parameters defined in unresolved foreground sky (Sec. 2.2.2)., and $C(\beta,N_{\nu})$ is a normalization factor. Because it is spectrally uncorrelated, pure thermal noise is irremovable but reducible with more integration time, while the $1/f$ noise is removable for data with high correlations between frequency channels. Olivari et al. (2016) explored the instrumental effects by adding thermal noise in the simulation. They presented the residual maps for the cases with no thermal noise and with thermal noise amplitudes equal to $0.05\mathrm{mK}$ and $0.08\mathrm{mK}$. It is concluded that the larger the thermal noise amplitude is, the less accurate is the reconstruction of the Hi power spectrum on small angular scales. A solution is also suggested by them in the case that we have a good estimate of the thermal noise power spectrum. The Hi + noise covariance matrix is used as the prior instead of the Hi covariance matrix, so that we recover the Hi signal plus thermal noise. We may correct for the thermal noise by using the estimate of the thermal noise power spectrum. For effects of the $1/f$ noise, readers may refer to Harper et al. (2018); Yohana et al. (2021). Therefore, we primarily focus on removing the foreground contamination, and will not discuss the impacts of thermal noise or $1/f$ instrumental noise in the subsequent discussion.
Table 1: Experimental specifications of simulations for the demonstration, the SKA-MID (Bacon et al., 2018; Wang et al., 2021) and the BINGO (Battye et al., 2012, 2013; Dickinson, 2014; Battye et al., 2016; Wuensche, 2019; Yohana et al., 2019; Abdalla et al., 2022; Wuensche et al., 2022) configurations, where the subscript FWHM stands for the full width of half maximum. Without loss of generality, we vary the channel width for convenience.
@@ 3 Foreground Removal Method
The Generalized Needlet Internal Linear Combination method was first applied to 21-cm intensity mapping data by Olivari et al. (2016). We will utilize this framework in general, but replace the theoretical cross-correlation matrix of the 21-cm signal with the one estimated by the robust principal component analysis method in one case. We will also elaborate on the refinement of the GNILC method and highlight the differences and improvements.
@@ 3.1 GNILC
We denote the observed sky at the pixel $p$ and frequency $i$ as $x_{i}(p)$, which sums the fluctuation of Hi  21-cm emission $s_{i}(p)$ and the foreground components $f_{i}(p)$. We remind the reader that $p$ does not necessarily mark the pixels in the real space, and it could be the index of the needlet coefficient in the GNILC. Similarly, the index $i$ may label any variant of frequency, such as the mode index of discrete cosine transform shown in Fig. 1. We may write the dependence of the frequency index in the vector form as
@@
where $\bm{x}(p)$, $\bm{s}(p)$, and $\bm{f}(p)$ are $n_{\text{ch}}$-dimensional column vectors and $n_{\text{ch}}$ is the number of frequency channels. The signal $\bm{s}$ is, of course, independent of the foreground $\bm{f}$. The bold characters denote the vectors (lowercase) or matrices (uppercase) defined in the frequency basis. From Eq. (19), the covariance matrix at pixel $p$ is
@@
to extract the Hi  signal $\hat{\bm{s}}$ (which is generally believed to deviate from $\bm{s}$) such that the total variance is minimized and the 21-cm signal is unbiased
@@
    where $\{\tilde{b}^{(j)}_{\ell}\}$ is the synthesis window family that ensures the reconstruction unbias condition $\sum_{j\in\mathcal{J}}\tilde{b}^{(j)}_{\ell}b^{(j)}_{\ell}=1$, $\Psi^{(j)}$ is the resultant smooth operator at needlet mode $j\in\mathcal{J}$. The symbol $\hat{}$ on $s$ means it is the estimate of $s$. At each frequency, we get the expected 21-cm map by summing up all needlet modes $\hat{s}(p)=\sum_{j\in\mathcal{J}}\Psi^{(j)}\hat{s}^{(j)}$.
@@ 3.1.1 Mixing matrix
The spectral window function determines how the needlet map is localized in both the spatial and spherical harmonics domains (refer to Appendix A). Thus the choice of the parameter $B$ should be in line with the spatial and spectral features of the 21-cm signal and foregrounds. Besides the dependence on the spectral window function of the needlet space, the determination of $\mathcal{D}$ is also affected by the foregrounds, masks, and telescope beams. We will show these effects in Sec. 4 and Sec. 5. In our analysis, we use the effective $\theta_{\mathrm{FWHM}}$ to quantify the needlet, beam, pixel size, and sky coverage in order to determine the domain $\mathcal{D}$ and $N_{p}$ the effective number of independent pixels in $\mathcal{D}$. First of all, bear in mind that we want a large $N_{p}$ to obtain an accurate estimate of the covariance matrix and reduce the bias (see Sec. 3.1.2). The independent size is limited by the map pixel size, the telescope beam, and the needlet size, which should be smaller than those three sizes. On the other hand, the maximum $\mathcal{D}$ is limited by the sky coverage (or mask), because if the $\mathcal{D}$ is significantly larger than the sky coverage, the data is averaged with the null field. In practice, the data and the null field are indeed blended together in the needlet domain, to a degree.
@@
where the superscript $(j)$ is omitted for simplicity. Note that any covariance matrix is a positive semi-definite matrix, and the signal is nonzero. Thus we can diagonalize the left-hand-side of Eq. (31) as
@@
So once we know the covariance matrix of the 21-cm signal, we can use Eq. (40) to obtain the mixing matrix. An essential precondition is to figure out the degrees of freedom of the foregrounds, i.e. $m$ values, which we will elaborate in more details in Sec. 3.1.2.
@@ 3.1.2 Degrees of freedom of the foregrounds
The above AIC only explicitly uses the likelihood of $\bm{x}(p)$, but in fact, the 21-cm prior maps can also be used when calculating the likelihood because they are the “prior knowledge” of reconstructed signal $\hat{\bm{s}}$. Here we do not distinguish the difference between the real signal and the prior because we assume their covariance matrices are very close. We will see from the following derivation, that this additional likelihood (except removable constants) is independent of the prior maps under certain assumptions.
@@
where $\hat{\bm{R}}_{\bm{s}}$, $\bm{R}_{\bm{s}}$ and $\bm{R}_{\bm{\delta}}$ are the covariance matrices of each term in Eq. (43). $\bm{C}_{\bm{s\delta}}$ is the covariance between the real 21-cm signal and the reconstruction error. We should notice that $\hat{\bm{R}}_{\bm{s}}$ is different from Eq. (39), because the latter is an intermediate estimation to calculate mixing matrix.
@@
where Eq. (50) is negligible because it is the second order of $\bm{\Delta}_{\bm{x}}$. The correlation $\bm{C}_{\bm{s\delta}}=\mathrm{E}(\bm{\delta}\bm{s}^{T})$ comprises correlations from Eqs. (47)-(49), and Eq. (47) will vanish because $\bm{s}$ and $\bm{f}$ are uncorrelated.
@@ 3.2 RPCA blind foreground removing
The noise effect is not tested in the RPCA-embedded GNILC approach. When extracting the Hi signals from the 21-cm IM data, the larger the thermal noise amplitude is, the less accurate is the reconstruction of the Hi power spectrum on small angular scales (Olivari et al., 2016). A solution is also suggested in the case that we have a good estimate of the thermal noise power spectrum. The Hi + noise covariance matrix is used as the prior instead of the Hi covariance matrix, so that we recover the Hi signal plus thermal noise. We may correct for the thermal noise by using the estimate of the instrumental noise power spectrum. This correction can also be applied to the RPCA-embedded GNILC approach where the prior covariance is replaced with the data covariance splitting. The sparse matrix comprises the Hi covariance and the thermal covariance (diagonal for simple models), which is roughly equivalent to the Hi plus noise prior. So we expect the performance of the RPCA-embedded GNILC approach is close to the GNILC with Hi signal plus thermal noise prior.

### v2

@@ Abstract
Intensity mapping (IM) of the neutral hydrogen 21-cm line has been studied as a promising technique to probe the large-scale structure (LSS) of the Universe (Madau et al., 1997; Battye et al., 2004; Masui et al., 2013; Santos et al., 2005; Bull et al., 2015; Santos et al., 2015; Bigot-Sazy et al., 2015, 2016; Li & Ma, 2017; Harper et al., 2018; Square Kilometre Array Cosmology Science Working Group et al., 2020; Xu et al., 2018; Anderson et al., 2018). Instead of measuring the emission from individual galaxies, IM measures integrated emission from a voxel of many galaxies in which the total signal can be detected by radio telescopes. Therefore, the integrated Hi  flux density in each pixel of sky maps contains multiple galaxies’ contributions to the signal.
Many experiments have been conducted or proposed to detect the Hi  power spectrum from auto- or cross-correlations. The techniques are mainly two kinds, the interferometer array and the single-dish (SD) measurement. The interferometer array measures the Fourier transform of the sky brightness in $uv$-plane (e.g. visibilities), which directly corresponds to the power spectra of the 21-cm signal. These include the Low-Frequency Array (LOFAR; Gehlot et al. 2019; Mertens et al. 2020), the Precision Array for Probing the Epoch of Reionization (PAPER; Parsons et al. 2010; Gogo et al. 2022), the Murchison Widefield Array (MWA; Trott et al. 2020), and the Hydrogen Epoch Reionization Array (HERA; DeBoer et al. 2017; Abdurashidova et al. 2022a, b; HERA Collaboration et al. 2023). The single-dish experiment, on the other hand, measures the specific flux of the 21-cm signal by direct imaging, which has achieved several detections in recent years via cross-correlation, including the HIPASS cross-correlation with 6dFGS galaxies (Pen et al., 2009), Green Bank Telescope (GBT) cross-correlation with DEEP2 galaxy survey (Chang et al., 2010) and with WiggleZ survey (Masui et al., 2013), Parkes telescope cross-correlation with 2dFGS galaxies (Anderson et al., 2018), and stacking of Parkes 21-cm fields with galaxy halos (Tramonte & Ma, 2020) and filaments (Tramonte et al., 2019). More recently, Cunnington et al. (2022) utilized $10.5$-hour observations of MeerKAT ($64$-dishes) drifted scan over $\sim 200\deg^{2}$ and cross-correlated it with WiggleZ galaxies, and achieved $7.7\sigma$ detection of the cross-correlation signal. In the near future, Five-hundred-meter Aperture Spherical radio Telescope (FAST, Nan et al. 2011; Bigot-Sazy et al. 2016), BINGO (Battye et al., 2012, 2013; Dickinson, 2014; Battye et al., 2016; Wuensche, 2019; Yohana et al., 2019; Abdalla et al., 2022; Wuensche et al., 2022), and SKA-MID (Square Kilometre Array Cosmology Science Working Group et al., 2020) will also utilize the SD IM technique to measure the 21-cm brightness of the sky and probe the LSS.
@@
This paper is organized as follows. In Sec. 2, we describe the models for the Hi  signal, foreground components, and instrumental noise. In Sec. 3, we present the mathematical framework of the eGNILC and make an extensive discussion of its function on foreground cleaning. Then in Sec. 4, we demonstrate the performance of the foreground removal methods, and apply it to the simulated data of the BINGO and SKA-MID configurations in Sec. 5. The conclusion is presented in Sec. 6.
@@ 2.1 Hi  21-cm intensity map
where in the last line, we utilized the fact that the Hi  signal traces the underlying matter density field on large scales after reionization. $\delta^{(3)}_{\mathrm{D}}$ is the 3-D Dirac delta function, $P_{\textsc{Hi}}\,$ and $P_{\mathrm{m}}$ are the Hi  and matter power spectra, $D(z)$ is the growth factor, and $b$ is the bias between the spatial distribution of the Hi  and the matter field. Then by ensemble-averaging the quadratic $a_{\ell m}$ (Eq. (6)), we obtain the angular power spectrum of 21-cm fluctuations in redshift bins ($z,z^{\prime}$; see also Battye et al. 2013)
@@ 2.2.2 Unresolved random sources
The spatial patterns are different in the DCT of power-law foregrounds because the synchrotron map is spatially smooth while temperature fluctuations are more dramatic in the free-free and point-source maps. The unresolved maps of both synchrotron and point sources are stochastic “noises”. Because the angular power spectrum of the unresolved synchrotron decreases faster ($\alpha$ is larger) than that of the unresolved point-source map, the map of the unresolved synchrotron foreground looks smoother than the unresolved point-source map. The horizontal stripes of the DCT images also suggest such spatial smoothness, while descending mode coefficients reflect the correlation of the $\nu$ and $\nu^{\prime}$ angular power spectrum. An outstanding feature of the Hi  signal is that the mode coefficients do not decay visibly.
@@ 2.2.3 Instrumental noise
The most common instrumental noise is the Gaussian distributed thermal noise. The rms of pixel noise is related to the system temperature, integration time, and frequency resolution (Bigot-Sazy et al., 2015; Olivari et al., 2016)
@@
where $f_{\rm k}$ is the knee frequency, $\alpha$ is the spectral index, $\omega$ is the inverse spectroscopic frequency wavenumber, $\Delta\nu$ is the total receiver bandwidth, $\beta$ is used to parametrize the spectral index of the PSD [^3] We remind the reader not to confuse the meaning of $\alpha$ and $\beta$ here with the parameters defined in unresolved foreground sky (Sec. 2.2.2)., and $C(\beta,N_{\nu})$ is a normalization factor. Because it is spectrally uncorrelated, pure thermal noise is irremovable but reducible with more integration time, while the $1/f$ noise is removable for data with high correlations between frequency channels. Olivari et al. (2016) explored the instrumental effects by adding thermal noise in the simulation. They presented the residual maps for the cases with no thermal noise and with thermal noise amplitudes equal to $0.05\mathrm{mK}$ and $0.08\mathrm{mK}$. It is concluded that the larger the thermal noise amplitude is, the less accurate the reconstruction of the Hi power spectrum on small angular scales is. They also suggest a solution in the case where there is a good estimate of the thermal noise power spectrum. The Hi +noise covariance matrix is used as the prior instead of the Hi covariance matrix, so that we recover the Hi signal plus thermal noise. We may correct for the thermal noise by using the estimate of the thermal noise power spectrum. For effects of the $1/f$ noise, readers may refer to Harper et al. (2018) and Yohana et al. (2021). Therefore, we primarily focus on removing the foreground contamination, and will not discuss the impacts of thermal noise or $1/f$ instrumental noise in the subsequent discussion.
Table 1: Experimental specifications of simulations for the demonstration, the SKA-MID (Square Kilometre Array Cosmology Science Working Group et al., 2020; Wang et al., 2021) and the BINGO (Battye et al., 2012, 2013; Dickinson, 2014; Battye et al., 2016; Wuensche, 2019; Yohana et al., 2019; Abdalla et al., 2022; Wuensche et al., 2022) configurations, where the subscript FWHM stands for the full width of half maximum. Without loss of generality, we vary the channel width for convenience.
@@ 3 Foreground Removal Method
The GNILC method was first applied to 21-cm IM data by Olivari et al. (2016). We will utilize this framework in general, but replace the theoretical cross-correlation matrix of the 21-cm signal with the one estimated by the RPCA method in one case. We will also elaborate on the refinement of the GNILC method and highlight the differences and improvements.
@@ 3.1 GNILC
We denote the observed sky at the pixel $p$ and frequency $i$ as $x_{i}(p)$, which sums the fluctuation of Hi  21-cm emission $s_{i}(p)$ and the foreground components $f_{i}(p)$. We remind the reader that $p$ does not necessarily mark the pixels in the real space, and it could be the index of the needlet coefficient in the GNILC. Similarly, the index $i$ may label any variant of frequency, such as the mode index of DCT shown in Fig. 1. We may write the dependence of the frequency index in the vector form as
@@
where $\bm{x}(p)$, $\bm{s}(p)$, and $\bm{f}(p)$ are $n_{\text{ch}}$-dimensional column vectors, and $n_{\text{ch}}$ is the number of frequency channels. The signal $\bm{s}$ is, of course, independent of the foreground $\bm{f}$. The bold characters denote the vectors (lowercase) or matrices (uppercase) defined in the frequency basis. From Eq. (19), the covariance matrix at pixel $p$ is
@@
to extract the Hi  signal $\hat{\bm{s}}$ (which is generally believed to deviate from $\bm{s}$) such that the total variance is minimized, and the 21-cm signal is unbiased
@@
    where $\{\tilde{b}^{(j)}_{\ell}\}$ is the synthesis window family that ensures the reconstruction unbias condition $\sum_{j\in\mathcal{J}}\tilde{b}^{(j)}_{\ell}b^{(j)}_{\ell}=1$, and $\Psi^{(j)}$ is the resultant smooth operator at needlet mode $j\in\mathcal{J}$. The symbol $\hat{}$ on $s$ refers to the estimate of $s$. At each frequency, we get the expected 21-cm map by summing up all needlet modes $\hat{s}(p)=\sum_{j\in\mathcal{J}}\Psi^{(j)}\hat{s}^{(j)}$.
@@ 3.1.1 Mixing Matrix
The spectral window function determines how the needlet map is localized in both the spatial and spherical harmonics domains (refer to Appendix A). Thus, the choice of the parameter $B$ should be in line with the spatial and spectral features of the 21-cm signal and foregrounds. In addition to the dependence on the spectral window function of the needlet space, the determination of $\mathcal{D}$ is also affected by the foregrounds, masks, and telescope beams. We will show these effects in Sections 4 and Sec. 5. In our analysis, we use the effective $\theta_{\mathrm{FWHM}}$ to quantify the needlet, beam, pixel size, and sky coverage in order to determine the domain $\mathcal{D}$ and $N_{p}$ the effective number of independent pixels in $\mathcal{D}$. First of all, bear in mind that we want a large $N_{p}$ to obtain an accurate estimate of the covariance matrix and reduce the bias (see Sec. 3.1.2). The independent size is limited by the map pixel size, the telescope beam, and the needlet size, which should be smaller than those three sizes. On the other hand, the maximum $\mathcal{D}$ is limited by the sky coverage (or mask), because if the $\mathcal{D}$ is significantly larger than the sky coverage, the data is averaged with the null field. In practice, the data and the null field are indeed blended together in the needlet domain, to a degree.
@@
where the superscript $(j)$ is omitted for simplicity. Note that any covariance matrix is a positive semidefinite matrix, and the signal is nonzero. Thus we can diagonalize the left-hand side of Eq. (31) as
@@
So once we know the covariance matrix of the 21-cm signal, we can use Eq. (40) to obtain the mixing matrix. An essential precondition is to figure out the degrees of freedom of the foregrounds, i.e. $m$-values, which we will elaborate in greater details in Sec. 3.1.2.
@@ 3.1.2 Degrees of freedom of the foregrounds
The above AIC only explicitly uses the likelihood of $\bm{x}(p)$, but in fact, the 21-cm prior maps can also be used when calculating the likelihood because they are the “prior knowledge” of reconstructed signal $\hat{\bm{s}}$. Here we do not distinguish the difference between the real signal and the prior because we assume their covariance matrices are very close. We will see from the following derivation that this additional likelihood (except removable constants) is independent of the prior maps under certain assumptions.
@@
where $\hat{\bm{R}}_{\bm{s}}$, $\bm{R}_{\bm{s}}$ and $\bm{R}_{\bm{\delta}}$ are the covariance matrices of each term in Eq. (43). $\bm{C}_{\bm{s\delta}}$ is the covariance between the real 21-cm signal and the reconstruction error. We should notice that $\hat{\bm{R}}_{\bm{s}}$ is different from that in Eq. (39), because the latter is an intermediate estimation to calculate mixing matrix.
@@
where Eq. (50) is negligible because it is the second order of $\bm{\Delta}_{\bm{x}}$. The correlation $\bm{C}_{\bm{s\delta}}=\mathrm{E}(\bm{\delta}\bm{s}^{T})$ comprises correlations from Eqs. (47) to (49), and Eq. (47) will vanish because $\bm{s}$ and $\bm{f}$ are uncorrelated.
@@ 3.2 RPCA blind foreground removing
The noise effect is not tested in the RPCA-embedded GNILC approach. When extracting the Hi signals from the 21-cm IM data, the larger the thermal noise amplitude is, the less accurate the reconstruction of the Hi power spectrum on small angular scales is (Olivari et al., 2016). A solution is also suggested in the case that we have a good estimate of the thermal noise power spectrum. The Hi + noise covariance matrix is used as the prior instead of the Hi covariance matrix, so that we recover the Hi signal plus thermal noise. We may correct for the thermal noise by using the estimate of the instrumental noise power spectrum. This correction can also be applied to the RPCA-embedded GNILC approach where the prior covariance is replaced with the data covariance splitting. The sparse matrix comprises the Hi covariance and the thermal covariance (diagonal for simple models), which is roughly equivalent to the Hi plus noise prior. So we expect the performance of the RPCA-embedded GNILC approach is close to the GNILC with Hi signal plus thermal noise prior.